There is monopolistic competition in the shampoo market. One of monopolistic firms

- DG. Company offers a special shampoo. Over a long period of time, G's total costs

(CK) of this special shampoo production (hi) are expressed by the following function

TC=3q3 - 27q? +100q. DG's total revenues (CZK) are given by the following quadratic

function TR = 63.25q - 6q?.

a) What amount of shampoo will DG produce?

b) What price will DG charge?

c) What profit will DG generate?

d) Present all results in an appropriate microeconomic model of G's behaviour.


1
Expert's answer
2021-11-28T17:48:39-0500

Solution:

a.). In monopolistic competition, the quantity produced will be where: MC = MR

Derive MC:

TC = 3q3 - 27q2 + 100q


MC = "\frac{\partial TC} {\partial Q}"= 9q2 – 54q + 100


MC = 9q2 – 54 + 100

 

Derive MR:

TR = 63.25q – 6q2

MR = "\frac{\partial TR} {\partial Q}" = 63.25 – 12q

 

MC = MR

9q2 – 54q + 100 = 63.25 – 12q

9q2 – 54q + 12q + 100 – 63.25 = 0

Q = 3.5

DG will produce 3.5 shampoos

 

b.). TR = P "\times" Q

63.25q – 6q2 = PQ

P = 63.25q – "\frac{6^{2} }{q}" = 63.25 – 6q

P = 63.25 – 6q = 63.25 – 6(3.5) = 63.25 – 21 = 42.25

Price = 42.25

DG will charge a price of 42.25

 

c.). Profit = TR – TC

Profit = (42.25 "\times" 3.5) – 3(3.53) – 27(3.52) + 100(3.5)

Profit = 147.875 – (128.625 – 330.75 + 350)

Profit = 147.875 – 147.875 = 0

DG will generate a profit of 0.